There are 5 blue socks, 3 red socs, and 2 green socks in a drwer. What is the probability of randomly choosing one blue sock, then one red soc, without putting the blue sock back first?

Answer by Alan3354(21583)

There are 5 blue socks, 3 red socs, and 2 green socks in a drwer. What is the probability of randomly choosing one blue sock, then one red soc, without putting the blue sock back first?
------------------
There are 5 blue sox out of a total of 10 sox, for the first choice, so the chance is 5/10 = 1/2.
For the 2nd pick, there are 3 red sox out of 9 total, so the chance is 3/9 or 1/3.
1/2 times 1/3 = 1/6